One year​ ago, your company purchased a machine used in manufacturing for $ 105 000. You have learned that a new machine is available that offers many​ advantages; you can purchase it for $ 155 000 today. It will be depreciated on a​ straight-line basis over ten years and has no salvage value. You expect that the new machine will produce a gross margin​ (revenues minus operating expenses other than​ depreciation) of $ 40 000 per year for the next ten years. The current machine is expected to produce a gross margin of $ 24 000 per year. The current machine is being depreciated on a​ straight-line basis over a useful life of 11​ years, and has no salvage​ value, so depreciation expense for the current machine is $ 9 545 per year. The market value today of the current machine is $ 65 000. Your​ company's tax rate is 42 %​, and the opportunity cost of capital for this type of equipment is 12 %. Should your company replace its​ year-old machine?
Answer #1

Solution:

Purchase price of Existing machine a $ 105,000
Depreciation per year b $ 9,545
Book value c=a-b $ 95,455
Market price of existing machine d $ 65,000
Capital loss e=c-d $ 30,455
Tax saving on capital loss f=e*42% $ 12,791.10
After tax Salvage value g=d+f $ 77,791.10
Machine 1 Machine 2
Cash outflow at t0 z -$ 77,791 -$ 1,55,000
Cash inflow per year h $ 24,000 $ 40,000
Depreciation per year i $ 9,545 $ 4,000
Cash inflow after depreciation j=h-i $ 14,455 $ 36,000
Cash flow after tax k=j*68% $ 9,829 $ 24,480
Add back depreciation l=k+i $ 19,374 $ 28,480
PVIFA @12%, 10 year m 5.6502 5.6502
PV of Cash inflow of 10 years n=l*m $ 109,470 $ 160,918
NPV n+z $ 31,679 $ 5,918

As continuation of existing machine gives more NPV than replacing the same with new one, hence the existing machine should be continued.